RC meeting and Run Plan, Monday October 31st
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Useful Links
Current Status
Polarized Target
- Current Polarization: N/A
- Requesting hourly screenshots of Target Strip Charts (and keep an eye on these during shift) - Will this be addressed during down??
- Experts performing a series of measurements this afternoon and tonight/overnight.
Run Status
- Small fire took beam down Friday ~1am; damage seen originally fixed around noon, but it was discovered that additional repairs needed.
- Transformer had to be replaced. Beam restored at 12:30am Saturday
- Good data taking Saturday Owl and Day, second half of Swing.
- A couple target motion resets were needed (Sat owl, Sat day, Sat swing..)
- Beam down 3-1/2 hours for tuning starting 15:30 on Saturday [Hall C beam halo was tripping Hall B DCs] - https://logbooks.jlab.org/entry/4072744
- Good data taking Sunday Owl, beam goes down 9am and is down through start of configuration change
- 9am - BPM readback failed.
- 12:45pm controlled access for software group to reboot/replace IOC
- 14:30pm begin crate replacement. Swapping power supplied fails, have to replace entire VME crate. Work stops shortly before midnight so that experts can get rest
- Decide to start configuration change 8am on Monday
Other Topics
- Hall in Controlled Access during morning, Power Permit for magnet-on target measurements. Will go back to CA later this afternoon: current estimate is around 4pm
- Plan on more measurements in power permit overnight (after oven has cooled down)
- Tomorrow morning, move to Restricted Access
- Shift Leaders on standby through end of Wednesday Day.
- Work in the hall so far:
- Radcon took swipes of target enclosure; will be ready to determine work controls for tomorrow
- Photodiode in target enclosure replaced
- Keith Cole and others continue work on replacing crate. Initial report was that the swap was done - still tests to do (DAQ test, possibly further work in the hall)
- Swapped the leads to the power supply of the 3rd Moller quad - https://logbooks.jlab.org/entry/4073979
- Xinzhan and Ben worked on SBS GEM issues - problems fixed? - https://logbooks.jlab.org/entry/4074106
- Other Subsystem updates?
Future Plans
- Still to do/schedule:
- Replace oven heater interlock box - Ellen
- Remaining BPM crate replacement work - Keith
- More SBS GEM issues??? (Xinzhan if there's anything left to do)
- Gas handling system - Carlos
- Any other access needed tonight, tomorrow, or Wednesday?
- MCC will set up separator on Wednesday (for 3-pass delivery)
- Most likely scenario is that we're ready to request beam sometime on Thursday day? Depends on timing of work in the hall.
- Long term "to do" items; not urgent, some can be done when 3He target not available
- (Luminosity scans with GEMs)
- (Opportunistic BCM calibration (higher current) before pass change?)
- (INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run) UPDATE: CAN BE DONE USING THE STAND-ALONE DAQ, SO NO LONGER ON THE OPPORTUNISTIC LIST.
- (Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29)
Current Run Plan
Configuration change.
Need updates for the 'standard' run plan below ------
NMR measurement
Restore beam (standard post-downtime procedure)
Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)
- Keep an eye out for drifting BPM positions
- Log all DAQ crashes in the excel worksheet along with the reason for the crash
- If beam delivery is interrupted for more than 4 hours, perform a checkout of the beam profile and raster follow Simona's procedure: [1]
- Cross check the rate with what is expected: [2]
- Cross check with the expected ion chamber readings: [3]
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Current Issues
- Shift Openings: Monday (morning) OWL, Tuesday OWL and SWING. [4]